OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4).

OpenStudy (mathlegend):

First, I would find the slope...

OpenStudy (mathlegend):

4-4/5-2 0/3 = 0

OpenStudy (mathlegend):

Then

OpenStudy (anonymous):

So the slope is 0??

OpenStudy (mathlegend):

Yes... it just means that the line would be horizontal.

OpenStudy (mathlegend):

y-y1=m(x-x1) y-4=0 +4 +4 --------- y = 4

OpenStudy (mathlegend):

@raerae01 do you understand?

OpenStudy (anonymous):

Yeah, thank you!